# DRPO: Efficient Reasoning via Decoupled Reward Policy Optimization

### v1

@@ 5.3 Length Reduction for Different Problem Difficulty
In this part, we study the impact of length reduction on questions with varying difficulties. In Figure 5(a), we present the performance of different methods across four math datasets of increasing difficulty, from GSM8K (easy) to AIME (very hard). As shown in Figure 5(a), we observe that: (1) on relatively easy questions like GSM8K, DRPO significantly reduces generation length from 1563 to 356 (77.2% length reduction) for 1.5B model and from 969 to 261 (73.1% length reduction) for 7B model with negligible performance drop (-1.1% for 1.5B model and -0.6% for 7B model). In contrast, all the baselines still bring dramatic performance degradation, indicated by a steep slope. (2) With question difficulty increasing from left to right in Figure 5(a), all the models exhibit longer reasoning length and introduce increasing performance drop, which is reasonable since difficult problems inherently require longer reasoning paths to solve. (3) When comparing Figure 5(a)(a) and (b), we can see that 7B models generally demand shorter reasoning length than 1.5B models and are more resilient to length reduction introduced by DRPO, as evidenced by flatter slopes than 1.5B models. (4) Across various difficulty levels, DRPO consistently achieves a better trade-off between performance and length than all the baselines, demonstrating its effectiveness in guiding efficient reasoning.
@@ 6 Conclusion and Discussions
A limitation of this study is that experiments were confined to 1.5B and 7B models due to computational constraints. Extending the proposed approach to larger models or broader reasoning tasks remains an interesting direction. Besides, while DRPO is designed for efficient reasoning, its formulation is general and can be extended to incorporate other rewards on positive data beyond length, such as process rewards. Another interesting direction is to incorporate mechanisms to adapt $\lambda$ to the difficulty level of questions, larger $\lambda$ for more difficult questions and smaller $\lambda$ for easier questions as implied by our results in Figure 5(a), which we leave for future investigation.
@@ A.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
Following prior work Rafailov et al. [2023], Go et al. [2023], we have
@@ A.2 Detailed Hyperparameter Setting
For all the methods, we employ the AdamW optimizer with a weight decay of 0.01 and set constant learning rate to $2\mathrm{e}^{-6}$ for 1.5B model and $1\mathrm{e}^{-6}$ for 7B model, following  Li et al. [2025a]. We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. The generation budget is limited to 8k tokens for both training and evaluation. The temperature is set to 0.6 for training.
@@
For the proposed DRPO, we follow  Li et al. [2025a] to set constraint value $\delta=1e^{-4}$, penalty constant $\beta_{0}=1e^{3}$, $\tau=10$. we tune regularization parameter $\lambda\in\{0.5,0.2,0.1\}$.
@@
RLOO-LP [Arora and Zanette, ] $r_{c}(o,q)-\alpha*r_{c}(o,q)*\sigma(\frac{|o|-\text{mean}\{|o_{i}|,r_{c}(o_{i},q)=1\}}{\text{std}\{|o_{i}|,r_{c}(o_{i},q)=1\}})$
ALP [Xiang et al., ] $r_{c}(o,q)-\beta*|o|*\max(\text{mean}\{r_{c}(o_{i},q)\},K^{-1})$
HAPO [Huang et al., ] $r_{c}(o,q)+w*\max\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),c\Big)r_{c}(o,q)+w*\min\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),0\Big)(1-r_{c}(o,q))$
L1-MAX [Aggarwal and Welleck, ] $r_{c}(o,q)*\text{clip}(\alpha(L_{T}-|o|)+\delta),0,1)$
SB [Yi et al., 2025] $\alpha*r_{c}(o,q)-\beta*\text{abs}\Big(|o|-L_{SOL}(q)\Big)$,
@@
LASER-D [Liu et al., 2025b] $r_{c}(o,q)+r_{c}(o,q)*\alpha\mathbb{I}(|o|\leq L_{A})$
@@ A.3 Limitation of incorporating length reward with group advantage
A.4 Detailed AES performance
In this part, we present detailed AES performance for each method in Table 4 and  5, where bold values denote the best AES performance for each method. We observe that all baseline methods yield negative AES scores for all settings, underscoring their inefficiency in preserving performance while reducing reasoning length. In contrast, DRPO consistently achieves positive AES scores for most cases, demonstrating its effectiveness in improving reasoning efficiency without sacrificing performance.
@@ A.4 Detailed AES performance
RLOO_LP ($\alpha$=0.1) 0.078 0.021 -0.423 -0.738 -0.172
RLOO_LP ($\alpha$=0.2) 0.246 0.325 -0.279 -0.694 -0.129
RLOO_LP ($\alpha$=0.4) 0.026 -0.13 -0.454 -1.18 -0.412

### v2

@@ 5.3 Length Reduction for Different Problem Difficulty
In this part, we study the impact of length reduction on questions with varying difficulties. In Figure 6(a), we present the performance of different methods across four math datasets of increasing difficulty, from GSM8K (easy) to AIME (very hard). As shown in Figure 6(a), we observe that: (1) on relatively easy questions like GSM8K, DRPO significantly reduces generation length from 1563 to 356 (77.2% length reduction) for 1.5B model and from 969 to 261 (73.1% length reduction) for 7B model with negligible performance drop (-1.1% for 1.5B model and -0.6% for 7B model). In contrast, all the baselines still bring dramatic performance degradation, indicated by a steep slope. (2) With question difficulty increasing from left to right in Figure 6(a), all the models exhibit longer reasoning length and introduce increasing performance drop, which is reasonable since difficult problems inherently require longer reasoning paths to solve. (3) When comparing Figure 6(a)(a) and (b), with two models from the same LLM family, we can see that 7B models generally demand shorter reasoning length than 1.5B models. Besides, models with larger sizes, e.g. 7B and 8B models, are more resilient to length reduction introduced by DRPO, as evidenced by flatter slopes than 1.5B models. (4) Across various difficulty levels, DRPO consistently achieves a better trade-off between performance and length than all the baselines, demonstrating its effectiveness in guiding efficient reasoning.
In Appendix A.5, we conducted an ablation study on the effect of different length reward designs in DRPO, comparing (1) a linear length reward, (2) a concave length reward and (3) cosine length reward. The results are summarized in Figure 7, which shows that the linear reward provides the broadest trade-off spectrum, while the concave and cosine rewards tend to yield higher accuracy at the cost of longer reasoning lengths.
@@ 6 Conclusion and Discussions
While DRPO is designed for efficient reasoning, its formulation is general and can be extended to incorporate other rewards on positive data beyond length, such as process rewards or other preference rewards. Extending the proposed approach to broader reasoning tasks remains an interesting direction. Another interesting direction is to incorporate mechanisms to adapt $\lambda$ to the difficulty level of questions, larger $\lambda$ for more difficult questions and smaller $\lambda$ for easier questions as implied by our results in Figure 6(a), which we leave for future investigation.
@@ A.1 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
Following prior work Rafailov et al. (2023); Go et al. (2023), we have
@@ A.2 Detailed Hyperparameter Setting
For all the methods, we employ the AdamW optimizer with a weight decay of 0.01 and set constant learning rate to $2\mathrm{e}^{-6}$ for 1.5B model and $1\mathrm{e}^{-6}$ for 7B model, following  Li et al. (2025a). We set the batch size to 128 for each step of RL, the mini-batch size to 32 for each iteration of model update, and sample 8 responses per question for training. The generation budget is limited to 8k tokens for both training and evaluation. The temperature is set to 0.6 for training.
@@
For the proposed DRPO, we follow  Li et al. (2025a) to set constraint value $\delta=1e^{-4}$, penalty constant $\beta_{0}=1e^{3}$, $\tau=10$. we tune regularization parameter $\lambda\in\{0.5,0.2,0.1\}$.
@@
RLOO-LP (Arora and Zanette, 2025) $r_{c}(o,q)-\alpha*r_{c}(o,q)*\sigma(\frac{|o|-\text{mean}\{|o_{i}|,r_{c}(o_{i},q)=1\}}{\text{std}\{|o_{i}|,r_{c}(o_{i},q)=1\}})$
ALP (Xiang et al., 2025) $r_{c}(o,q)-\beta*|o|*\max(\text{mean}\{r_{c}(o_{i},q)\},K^{-1})$
HAPO (Huang et al., 2025) $r_{c}(o,q)+w*\max\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),c\Big)r_{c}(o,q)+w*\min\Big(\cos(\min(\frac{\pi}{2}\frac{|o|}{h(q)},\pi)),0\Big)(1-r_{c}(o,q))$
L1-MAX (Aggarwal and Welleck, 2025) $r_{c}(o,q)*\text{clip}(\alpha(L_{T}-|o|)+\delta),0,1)$
SB (Yi et al., 2025) $\alpha*r_{c}(o,q)-\beta*\text{abs}\Big(|o|-L_{SOL}(q)\Big)$,
@@
LASER-D (Liu et al., 2025b) $r_{c}(o,q)+r_{c}(o,q)*\alpha\mathbb{I}(|o|\leq L_{A})$
@@ A.3 Limitation of incorporating length reward with group advantage
A.4 Experiments on Non-mathematical Reasoning Task
To evaluate DRPO’s generalization on non-mathematical reasoning tasks, we conducted additional experiments on the logic puzzle reasoning task. Following Xie et al. (2025); Su et al. (2025), the training dataset is limited to 3 to 7-person logic puzzles with K&K logic puzzle dataset (Xie et al., 2024) and the test dataset contains 2 to 8-person puzzles. We train 1.5B models for all methods for 400 steps and conduct evaluation every 80 steps. As shown in Figure 6, DRPO method still exhibits a substantially better trade-off than all other baselines, demonstrating the generalizability of the proposed method to other tasks. Notably, DRPO significantly reduces the generation length from 2095 to 1400 (33.2% length reduction) without any performance drop (from an accuracy of 0.972 to 0.974 ).
Figure 6: Comparison of performance-efficiency trade-off on logical puzzle reasoning task. Squares denote models trained with reference methods without length penalties, i.e., $\lambda$$\infty$ (corresponding to DisCO) for DRPO, $\alpha=0$ for RLOO-LP, $\beta=0$ (corresponding to GRPO) for ALP, $w=0$ for HAPO. Triangles denote the models trained by other works.
A.5 Ablation study on length reward design
[…]
Figure 7: Ablation studies on different reward designs. Left is the length reward values with respect to $|o|/C$; Right is the performance-efficiency trade-offs for different reward designs on mathematical reasoning task. Square denotes the model trained without length penalties, i.e., $\lambda$$\infty$ (DisCO).
A.6 Detailed AES performance
In this part, we present detailed AES performance for each method in Table 4,  5, and  6, where bold values denote the best AES performance for each method. We observe that all baseline methods yield negative AES scores for almost all settings, underscoring their inefficiency in preserving performance while reducing reasoning length. In contrast, DRPO consistently achieves positive AES scores for most cases, demonstrating its effectiveness in improving reasoning efficiency without sacrificing performance.
@@ A.6 Detailed AES performance
RLOO-LP ($\alpha$=0.1) 0.078 0.021 -0.423 -0.738 -0.172
RLOO-LP ($\alpha$=0.2) 0.246 0.325 -0.279 -0.694 -0.129
RLOO-LP ($\alpha$=0.4) 0.026 -0.13 -0.454 -1.18 -0.412
@@
Table 6: Detailed AES performance for 8B models.
Method GSM8K MATH500 OlympiadBench AIME Ave.
RLOO-LP ($\alpha$=0.1) 0.378 0.198 0.282 0.295 0.212
RLOO-LP ($\alpha$=0.2) 0.515 0.203 0.313 0.315 0.251
RLOO-LP ($\alpha$=0.4) 0.064 -0.126 0.127 0.392 -0.119
ALP ($\beta$=1e-9) 0.476 0.215 0.053 -0.67 -0.01
ALP $\beta$=1e-8) 0.22 -0.252 -0.686 -2.405 -0.584
ALP ($\beta$=1e-7) -2.139 -1.773 -1.81 -3.733 -2.289
HAPO (ww=0.01) -0.021 -0.17 -0.251 -0.755 -0.413
HAPO (ww=0.1) 0.197 -0.264 -0.445 -0.673 -0.366
HAPO (ww=1) -0.23 -0.667 -0.542 -1.579 -0.775
DRPO ($\lambda$=0.5) 0.29 0.172 0.098 0.091 0.122
DRPO ($\lambda$=0.2) 0.607 0.435 0.313 0.068 0.296
DRPO ($\lambda$=0.1) 0.624 0.449 0.292 0.04 0.297
@@ A.7 Case Study
LLM usage
We use LLM only to help correct grammar errors and polish writing.